OpenStudy (anonymous):

p=2L+2W L=96 and w=144 i think that p=480 is that right?

OpenStudy (lgbasallote):

P = 2(96) + 2(144) P = 192 + 288 P = 480 right

OpenStudy (anonymous):

OK thank you